OpenStudy (babyslapmafro):

Please help me find the Laplace transform of f(t)=sin(5t)cos(4t)

OpenStudy (holsteremission):

Try rewriting as \(\sin(5t)\cos(4t)=\dfrac{\sin t+\sin(9t)}{2}\). If you know the transform for \(\sin(at)\), you're as good as done.

OpenStudy (babyslapmafro):

Ok I'll try that

OpenStudy (babyslapmafro):

Thanks